OpenStudy (chloelourdes):

chris and his brother leave the same point and run in opposite directions onna straight road. after running for 3 hours, they are 18 miles apart. if chris runs an average of 1mph faster than his brother, what is chris's average speed?

OpenStudy (will.h):

1st you want to find the velocity of them both when they 1st kick off \[v = \frac{ d }{ t }\] where d is distance and t is time \[v = \frac{ 18 }{ 3 }\] \[v = 6\] that's the speed but then if chris runs 1mph faster than his brother then you should add 1mph to that which means 6 + 1 = 7mph is the average velocity of chris